OpenStudy (anonymous):

how do you reduce -5/8 / 11/4 would i just divide it?

OpenStudy (angela210793):

a/b : c/d=a/b*d/c follow this rule

OpenStudy (anonymous):

cross multiply.

OpenStudy (anonymous):

-5/8 * 4/11

OpenStudy (anonymous):

but it's a fraction over a fraction. ohh so i flip the second fraction?

OpenStudy (angela210793):

yea tht's wht u do :)

OpenStudy (anonymous):

ok thanks!

OpenStudy (phi):

Does this make sense? Multiply top and bottom by 4/11 to make the messy denominator become 1 \[\frac{\frac{-5}{8}} {\frac{11}{4}}=\frac{\frac{-5}{8}\cdot\frac{4}{11}} {\frac{11}{4}\cdot\frac{4}{11}}=\frac{\frac{-5}{8}\cdot\frac{4}{11}}{1} =\frac{-5}{8}\cdot\frac{4}{11}\]